# Annotation Free Semantic Segmentation with Vision Foundation Models

### v1

@@ 3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of possible categories $\{g_{1},g_{2},\dots,g_{L}\}$. A transformer-based image encoder [12] takes as input an image ${im_{i}}\in\mathbb{R}^{C,H,W}$, divided into patches $im_{i}=[\mathbf{im}^{1}_{i},\dots\mathbf{im}^{p}_{i},\dots,\mathbf{im}^{N}_{i}]$ where $N$ is the number of patches. It then extracts a class token and patches embedding $x_{i}=[\mathbf{cls}_{i},\mathbf{x}^{1}_{i},\dots,\mathbf{x}^{p}_{i},\dots,\mathbf{x}^{N}_{i}]$ where $\mathbf{cls}_{i},\mathbf{x}^{p}_{i}\in\mathbb{R}^{D}_{v}$. CLIP [26] provides an image and a text encoder. The text encoder takes as input a text description, e.g. {a photo of a “class”}, and produces a text feature $\mathbf{t}_{i}$ followed by linear projection to $\mathbb{R}^{D}$. Another linear projection layer maps $\mathbf{cls}_{i}$ from the vision encoder to a joint space with the text features $\mathbb{R}^{D}$. CLIP is trained to maximize the cosine similarity between the image class token $\mathbf{cls}_{i}$ and the text representation $\mathbf{t}_{i}$. The patch features $\{\mathbf{x}^{p}_{i}\}$ can be similarly projected onto the same space $\mathbb{R}^{D}$, however we found that the linear projection layer applied directly to the patches features results in features that are sub-optimally aligned with corresponding text features. To overcome this, we follow a procedure similar to Zhou et al. (2022)  [43], where we apply the value projection $V$ of the self attention layer to the input features of the last transformer block, followed by a layer norm, ignoring the key and query. Refer to the Appendix for more details.
3.2 Stage.1: Object detection & masks generation
In this section, we outline our strategy to generate pseudo semantic segmentation labels using  CLIP and SAM. We employ CLIP to recognize categories present in the image. Stage 1.1 extracts query points for SAM mask generation while Stage 1.2 complements this step by labeling SAM’s automatically generated masks. We first describe the detection of categories present in an image. Then, we detail our pseudo mask generation.
High-resolution feature extraction We divide the image it into $C$ crops in a sliding window fashion, process each crop separately with CLIP and rearrange the patches returned from CLIP to construct the features for the full image. This guarantees precise and high quality feature maps for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$. We refer to the Appendix for more details.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop to a set of text features $T$ of a vocabulary. Specifically, all possible labels in a given dataset are considered as the vocabulary, e.g., for COCO-Stuff a set of 171 classes.
@@
Object Detection. Each crop is classified with an object category when the object’s text feature $\textbf{t}_{o}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all text features T in the vocabulary. An object is considered as present in the image if it has been assigned to more than a predefined number of crops (see appendix for more details).
Pseudo mask generation. We compute a similarity matrix between the full image feature map and the text features for the detected categories. For each category we select 5 patches with the highest similarity to their corresponding text feature as query points. We feed these query points, along with the original image to SAM and select the mask with the highest confidence (Fig 2).
@@
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig 5). Therefore, stage 1.2 performs a complementary pseudo label generation using CLIP and SAM to further boost the performance.
@@
Automatic mask generation. As an alternative to stage 1.1, we rely on all possible masks extracted from the full image using SAM’s automatic mask generation pipeline. We constrain the masks by size and predicted IOU to filter out the low quality and duplicate masks (See appendix).
@@
Mask labelling In order to classify each generated mask by SAM, we compute CLIP’s mean feature embedding corresponding to the area covered by the mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask. We refer to the supplementary for comparison of each step results.
Figure 2: Query point selection and mask generation with SAM: Our method selects 5 patches with the highest similarity to the detected object’s text embedding. The coordinates for the center of these patches are forwarded to SAM which generates 3 different segmentation masks for the object. Our method selects the segmentation mask with the highest confidence.
3.3 Stage.2: Lightweight semantic segmentation
@@
The method, as described so far, provides a way to generate fine semantic segmentations from an image dataset with a predefined vocabulary. The generated masks do not cover every object in an image and might be wrong sometimes. It can be also rendered inefficient when constrained with low compute, as it involves querying two foundation models. To overcome this, we are now concerned with how to align any off-the-shelf pretrained vision encoder with text semantics, and with no human supervision.
@@
Alignment Module. We consider the predicted masks along with their corresponding predicted categories, as produced by our method in 3.2, as pseudo labels which we use to train a simple alignment module to map the image patch features to the text embedding. Such a mapping would ground any pretrained vision encoder with language, for dense prediction tasks. Since supervised models may show a bias towards some object categories, we focus on self-supervised pretrained models. Namely, we consider DINOv2 [25], a recent model trained in a fully self-supervised fashion with no text alignment. To overcome the noisy pseudo annotations we rely on 1) frozen pretrained text features as anchors, 2) already discriminative image patches features 3) a loss function that is robust to noise.
Pseudo label generation. Let $\mathbf{D}$ be a dataset of unlabelled images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$. Using the first stage of the pipeline, we extract object masks for each image $im_{i}$ along with their assigned text features $\{m_{i}^{s},t_{s}\}$. The output of the image encoder (e.g. DINOv2) is $\mathbf{x}_{i}=[\mathbf{x}^{1}_{i},\dots,\mathbf{x}^{p}_{i},\dots,\mathbf{x}^{N}_{i}]$ where $N$ is the number of patches, ignoring the $\mathbf{cls}$ token. Using the generated masks and their associated detected categories we assign to each patch $\mathbf{x}^{p}_{i}$ a pseudo label $y^{p}_{i}\in\{1,..,K\}$ where $K$ is the number of detected categories in the dataset $\mathbf{D}$. From the text encoder we extract $K$ text features $T=\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{K}\}$.
@@
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [1, 44] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
@@
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [16] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
@@
Inspired by supervised contrastive loss SupCon [18] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
@@
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [18] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [2] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs, as we show in the loss ablation, Sec 4.4.1. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs.
Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder. To provide pixel-level predictions we interpolate the similarities to the original image dimension.
@@ 4 Experiments
We evaluate our approach on semantic segmentation datasets followed by ablation of our model and loss design.
@@ 4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. Refer to the Appendix for a detailed description. We use COCO-Stuff [5] images for pretraining our alignment model, composed of only 118k images. We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline, Stage.1 3.2. Then, we train our alignment module on the pseudo annotations for 10 epochs.
Figure 3: Patch-level alignment between image and class. First row shows the original image from Pascal VOC. Second row shows the similarity between the high-resolution patch features from CLIP-L/14 and the image-level label. Third row shows the similarity map after aligning a DINOv2-L/14 model with FMbSeg.
Figure 4: Qualitative results of zero-shot segmentation. We show 2 samples for each one of the considered datasets. The first row shows the ground truth labels. The second row shows the results of FMbSeg.
Figure 5: Qualitative evaluation of Stage.1.1. SAM query points are shown in green stars. Top sub-figure shows instances of correct detected labels and segmentations by Stage.1.1. Below demonstrates its limitations. Small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels all image masks generated by SAM.
Datasets. Pascal VOC [13] contains 20 foreground classes and everything else is labeled as a background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and PV-21 with the background class. Pascal Context [24] (PC-59) contains 59 classes of objects and stuff. COCO Stuff [5] (CS-171) contains 80 different objects categories and 91 stuff classes. We evaluate on both object categories only (CO- 80) and both objects and stuff categories (CS- 171). We further consider CityScape [9] (City) and ADE20K [42] (ADE). For PV-21, we add different Stuff categories to predict the background class.
@@
GroupVit [34] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [43] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [29] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [7] 15M + Labels. 55.0 83.2 33.9 31.6 22.4 24.0 17.1
CLIP-S4 [16] 0.12M - - 33.6 - 22.1 - -
SCCLIP [31] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [33] - 59.0 - 30.4 - - - -
CaR [30] - 67.6 - 30.5 36.6 - -
SAM-CLIP [32] 41M 60.6 - 29.2 31.5 - 17.1
FMbSeg (ours) 0.12M 67.73 85.65 42.72 57.63 29.88 28.36 16.25
@@
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require.
Image Level Annotation: GroupVit [34] and TCL [7]. Annotation free methods: We consider Mask CLIP [43] (the best performing variant), CLIP-S4 [16] and SAM-CLIP [32]. Training free methods: CLIP-DIY [33], SCCLIP [31], CaR [30] and ReCo [29].
@@
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [7]. We don’t apply any post-refinement and only one standard template for text features is considered unlike other approaches using 80 different templates.
4.2 Annotation Free Semantic Segmentation Results

### v2

@@ 3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of categories in the vocabulary $V=\{class_{1},\dots,class_{L}\}$. CLIP’s transformer-based image encoder [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.] takes as input an image ${\mathbf{im}_{i}}\in\mathbb{R}^{C,H,W}$ divided into patches $\mathbf{im}_{i}=[im^{1}_{i},\dots im^{p}_{i},\dots,im^{N}_{i}]$ and extracts a class token $\mathbf{cls}_{i}$ and patches embedding $\mathbf{x}_{i}=[x^{1}_{i},\dots,x^{p}_{i},\dots,x^{N}_{i}]$. CLIP’s text encoder takes as input a text description {a photo of a “$class_{i}$”} and produces the text feature $\mathbf{t}_{i}$, with $class_{i}$ $\in V$ corresponding to the image label. Visual and textual features represented by $\mathbf{cls}_{i}$ and $\mathbf{t}_{i}$ are separately projected to a joint embedding space $\mathbb{R}^{D}$ and the cosine similarity between them is maximized during CLIP’s training. While the original CLIP architecture discards the patch features $\mathbf{x}_{i}$, they can be projected onto the same space $\mathbb{R}^{D}$. This would enable us to compute the similarity of any category in $V$ with individual patches $\mathbf{x}^{p}_{i}$ and produce a rough localization of the objects in the image. Refer to the Appendix for more details on the specific architecture for projecting patch features $\mathbf{x}_{i}$ to $\mathbb{R}^{D}$.
3.2 Stage 1: Object detection & masks generation
In this section, we outline our strategy to generate pseudo semantic segmentation labels using  CLIP and SAM. We employ CLIP to recognize categories present in the image and SAM for mask generation. We propose two complementary methods for this, Stage 1.1 and Stage 1.2. We provide further details and examples in the Appendix.
@@ 3.2 Stage 1: Object detection & masks generation
High-resolution feature extraction We (over-)sample each image into a high resolution, divide it into $C$ crops in a sliding window fashion, process each crop separately with CLIP and rearrange the patches returned from CLIP (section 3.1) for all crops to construct the features for the full image. This guarantees precise and high quality feature maps for each image. Besides, for each crop $c$, we extract a classification token $\mathbf{cls}_{c}$.
Defining the set of concerned categories. To detect classes present in an image our method computes the similarity between the classification token for each crop to a set of text features corresponding to descriptions extracted from vocabulary $V$ (section 3.1). All possible labels in a given dataset are considered as the vocabulary (e.g. a set of 171 classes for COCO-Stuff).
Object presence detection. Each crop $c$ is classified with an object category when the object’s text feature $\textbf{t}_{o}$ has the highest similarity to the classification token $\mathbf{cls}_{c}$ among all descriptions extracted from vocabulary $V$. An object is considered as present in the image if it has been assigned to more than a predefined number of crops (set to 1 in our experiments).
Pseudo mask generation. We compute a similarity matrix between the full image feature map and the text features for the detected categories (Fig. 2). For each category we select 5 patches with the highest similarity as query points. We feed these points along with the original image to SAM and select the mask with the highest confidence.
@@
Stage 1.1 may ignore small objects or generate partial masks for an image (Fig. 3) due to sub-optimal query points. Stage 1.2 performs a complementary pseudo label generation using CLIP and SAM to further boost the performance. We refer to the supplementary for comparison of the performance for stage 1.1 and stage 1.2.
@@
Automatic mask generation. We retrieve all possible masks extracted from the full image using SAM’s automatic mask generation pipeline. We constrain the masks by size and predicted IOU to filter out the low quality and duplicate masks (See appendix).
@@
Mask labelling In order to classify each generated mask by SAM, we compute CLIP’s mean feature embedding corresponding to the area covered by the mask and compute its similarity to the text features of the detected categories in the image. The class with the highest similarity is selected as the pseudo label for the corresponding mask.
3.3 Stage 2: Lightweight semantic segmentation
@@
Stage 1 extracts segmentation masks from a dataset with a predefined vocabulary. However, the generated masks might be noisy. Additionally, it involves querying two foundation models and might render inefficient when constrained with low compute. To overcome this, we propose to use annotations generated in Stage 1 to align any off-the-shelf pretrained vision encoder with text semantics and with no human supervision.
@@
Alignment module. We consider the generated masks in 3.2 as pseudo labels for training a simple alignment module that maps the image patch features to the text embedding. Such a mapping would ground any pretrained vision encoder with language for dense prediction tasks. Since supervised models may show a bias towards some object categories, we focus on self-supervised pretrained models. Namely, we consider DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], a recent model trained in a fully self-supervised fashion with no text alignment. To overcome the noisy pseudo annotations we rely on 1) frozen pretrained text features as anchors, 2) already discriminative image patches features 3) a loss function that is robust to noise.
[…]
@@
Training the alignment module. The pretrained image encoder remains frozen and we optimize a small alignment module $\mathcal{M}$ to map the patch representations $\{\mathbf{x}^{p}_{i}\}$ to the CLIP text embedding space $z^{p}_{i}=\mathcal{M}(\mathbf{x}^{p}_{i})\in\mathbb{R}^{D}$. Note that the method is agnostic to the specific encoder used. As we strive for simplicity, we design our alignment module as one transformer block with multi-head self-attention layer. The self-attention layer allows each patch to attend to other patches in the image. Since we aim for an open-set semantic segmentation, cross-attention with text features [Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al., Zou et al.(2023)Zou, Dou, Yang, Gan, Li, Li, Dai, Behl, Wang, Yuan, et al.] is not used as it would require a joint processing of the image features and a closed set of text features at test time.
@@
For notation clarity, we drop the image index and consider only an across-batch patch index $i;i\in\{1,\dots,N*B\}$ where N is the number of patches in an image and $B$ is the batch size. CLIP [He et al.(2023)He, Jamonnak, Gou, and Ren] contrasts the similarity of text features with image class tokens using a cross entropy loss, where a one to one correspondence exists between an image class token and its text features. In our case, we have many image features $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$ extracted from patches of many images and few text features corresponding to detected categories in $\mathbf{D}$. Upon early experiments with CLIP loss, we found it not scalable to image patches and exhibiting poor convergence.
@@
Inspired by supervised contrastive loss SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] that operates on positive and negative pairs, we construct two types of pairs: pairs of image patches (patch-patch pairs: $<\mathbf{z}_{i},\mathbf{z}_{j}>$) and pairs of image patches and text features (patch-text pairs: $<\mathbf{z}_{i},\mathbf{t}_{k}>$). A patch-patch pair $<\mathbf{z}_{i},\mathbf{z}_{j}>$ is considered positive if the patches belong to the same category class $y_{i}=y_{j}$ and negative otherwise. Patch-text pairs $<\mathbf{z}_{i},\mathbf{t}_{k}>$ are positive if $y_{i}=k$. We construct a loss function of two terms operating on the two types of said pairs:
@@
where $N_{y_{i}}$ is the number of patches in the batch that have the same label as $y_{i}$. This loss term operates on the image patches and it maximizes the similarities of the patches of the same category across different images. This term is a generalization of SupCon [Khosla et al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola, Maschinot, Liu, and Krishnan] to image patches with no data augmentation.
A similar loss function has been proposed for image classification in [Aljundi et al.(2023)Aljundi, Patel, Sulc, Chumerin, and Reino] and it was shown that $\ell_{t}$ (3) formulation can be expressed as a smooth approximation to the maximum function of SupCon and Cross Entropy (CE) loss; in our case of SupCon on patch-text pair and CE on the patch feature with the corresponding text feature as a representative class prototype. This approximation is key to allow a smooth optimization of the different similarities while tolerating possible noisy pairs. We optimize the alignment module $\mathcal{M}$ with $\ell_{\text{TSupCon}}$ (Eq.  1), for a fixed number of epochs and provide ablation on different loss functions in the Appendix. Deployment. At test time, we extract $\{\mathbf{x}_{i}\}$ features from the image, using the frozen image encoder. The image patches are forwarded through $\mathcal{M}$, $\mathbf{z}_{i}=\mathcal{M}(\mathbf{x}_{i})$, then we compute for each $z_{i}$ the most similar text feature $\mathbf{t}_{k}$, after which we assign to $\mathbf{z}_{i}$ a label $k$. Text features are precomputed using a frozen text encoder. To provide pixel level predictions we interpolate the similarities to the original image dimension.
@@ 4.1 Experimental Setup
Implementation Details. We consider Vit-L/14 for both CLIP and DINOv2. We train our alignment module with SGD optimizer, cosine annealing scheduler and a batch size of 5 images. We use COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] images for pretraining our alignment model, composed of only 118k images. We extract pseudo segmentations of COCO-Stuff vocabulary using the first stage of our pipeline (section  3.2). Next, we use them to train our alignment module for 10 epochs. Since Stage 1.2 generates many more masks per image compared to Stage 1.1, in order to balance the pseudo annotations of these complementary methods, we randomly sample $75\%$ of annotations from stage 1.2 (see Appendix for ablation).
Figure 2: Patch level alignment between image and class. First row shows images from Pascal VOC. Second row shows the similarity between patch features from CLIP and the text features of the detected category. Third row shows the similarity map after aligning a DINOv2 model with FMbSeg.
Figure 3: Qualitative evaluation of Stage.1.1. SAM query points generated by our method are shown in green stars. Left shows instances of correct segmentations by Stage.1.1. and Right demonstrates its limitations; small objects, wrongly detected classes (due to ambiguities) and not enough query points to cover all instances. Stage 1.2 alleviates the issue with small objects and incomplete masks since it labels the masks generated accurately by SAM.
Figure 4: Qualitative results of zeroshot segmentation. The first row shows the ground truth labels. The second row shows the results of FMbSeg.
Datasets. COCO Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] contains 80 things categories and 91 stuff classes. We evaluate on both things only (CO-80) and things + stuff (CS-171) categories. Pascal VOC [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman] contains 20 foreground classes and everything else is labeled as background. To isolate the effect of the background class we evaluate on both (PV-20) excluding the background class and (PV-21) with the background class. For PV-21, we add stuff classes from CS-171 as background categories. Pascal Context [Mottaghi et al.(2014)Mottaghi, Chen, Liu, Cho, Lee, Fidler, Urtasun, and Yuille] (PC-59) contains 59 classes of objects and stuff. CityScapes  [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] (City) contains 30 categories from street view images. ADE20K [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] (ADE) contains 150 categories from various indoor and outdoor scenes.
@@
GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] 41M +Labels 50.4 79.7 18.7 27.5 15.3 11.1 9.2
Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] - 38.8 74.9 23.6 20.6 16.4 12.6 9.8
ReCo [Shin et al.(2022)Shin, Xie, and Albanie] - 25.1 57.7 19.9 31.6 14.8 21.1 11.2
TCL [Cha et al.(2023)Cha, Mun, and Roh] 15M + Labels. 55.0 83.2 33.9 31.6 22.4 24.0 17.1
CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] 0.12M - - 33.6 - 22.1 - -
SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille] - 59.1 - 30.4 - 22.4 - -
CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] - 59.0 - 30.4 - - - -
CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] - 67.6 - 30.5 36.6 - -
SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] 41M 60.6 - 29.2 31.5 - 17.1
FMbSeg 0.12M 67.73 85.65 42.72 57.63 29.88 28.37 16.25
@@
Compared Methods. We compare with methods that target the alignment of image features with language for open vocabulary semantic segmentation. We divide methods according to the annotation and training they require. Image Level Annotation: GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and TCL [Cha et al.(2023)Cha, Mun, and Roh]. Annotation-free methods: We consider Mask CLIP [Zhou et al.(2022)Zhou, Loy, and Dai] (the best performing variant), CLIP-S4 [He et al.(2023)He, Jamonnak, Gou, and Ren] and SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari]. Training free methods: CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni], SCCLIP [Wang et al.(2023a)Wang, Mei, and Yuille], CaR [Sun et al.(2023)Sun, Li, Torr, Gu, and Li] and ReCo [Shin et al.(2022)Shin, Xie, and Albanie].
@@
Metrics. We consider the widely adopted Mean Intersection over Union (mIoU). We follow the evaluation protocol of TCL [Cha et al.(2023)Cha, Mun, and Roh]. We don’t apply any post-refinement and only one standard template {a photo of a “$class$”} for text descriptions is considered unlike other approaches using 80 different templates during evaluation.
4.2 Annotation-Free Semantic Segmentation Results